# NRGPT: An Energy-based Alternative for GPT

### v1

@@ F.3 Listops
We perform experiment on nested math operations on lists of integers, which are a version of ListOps (nangia2018listops). Our ListOps setting consists of three functions: maximum, median and sum modulo 20. Our inputs range from 0 to 19.
@@ F.5 OpenWebText
We perform experiments on natural language modeling using the OpenWebText corpus, which is an open-source recreation of GPT-2’s WebText dataset (radford2019language). The dataset contains approximately 17GB of text with 9B tokens that came from 8 million documents. We tokenize using byte-pair encoding (BPE) with a vocabulary size of 50,257 tokens. Our training sequences are fixed-length contexts of 1024 tokens. Table˜7 lists the hyperparameters and respective values/ranges used in our experiments.
@@ F.5.1 MMLU
Perplexity and simple generation metrics are insufficient to understand the performance of language models on downstream tasks. Thus, we additionally characterize the equal-parameter models of Table˜5 on the MMLU dataset (hendrycks2020measuring) using five-shot generation in Table˜6. Remarkably, we find that NRGPT outperforms both GPT and RecGPT at constant parameter count on the average MMLU score.
As a caveat, we note that transformers at the ${\sim}125$M scale generally perform poorly on these downstream tasks. For example, the original MMLU paper (hendrycks2020measuring) shows that GPT-3-small (2.7 billion params) and GPT-3-medium (6.7 billion params) (brown2020language) achieve a 25% success rate — the exact same as random guessing. A large GPT-2 model (1.5 billion params) (radford2018improving) achieves a slightly better 32.4% accuracy, but this is still only 7% higher than random prediction. Our NRGPT model at 125M parameters (an order of magnitude smaller than the MMLU results of GPT-2) achieves a 29.3% accuracy, which is the best performance per parameter count on this task that we are aware of.

### v2

@@ F.3 Listops
We perform experiment on nested math operations on lists of integers, which are a version of ListOps [Nangia and Bowman, 2018]. Our ListOps setting consists of three functions: maximum, median and sum modulo 20. Our inputs range from 0 to 19.
@@ F.5 OpenWebText
We perform experiments on natural language modeling using the OpenWebText corpus, which is an open-source recreation of GPT-2’s WebText dataset [Radford et al., 2019]. The dataset contains approximately 17GB of text with 9B tokens that came from 8 million documents. We tokenize using byte-pair encoding (BPE) with a vocabulary size of 50,257 tokens. Our training sequences are fixed-length contexts of 1024 tokens. Table˜7 lists the hyperparameters and respective values/ranges used in our experiments.
@@ F.5.1 MMLU
Perplexity and simple generation metrics are insufficient to understand the performance of language models on downstream tasks. Thus, we additionally characterize the equal-parameter models of Table˜5 on the MMLU dataset [Hendrycks et al., 2020] using five-shot generation in Table˜6. Remarkably, we find that NRGPT outperforms both GPT and RecGPT at constant parameter count on the average MMLU score.
As a caveat, we note that transformers at the ${\sim}125$M scale generally perform poorly on these downstream tasks. For example, the original MMLU paper [Hendrycks et al., 2020] shows that GPT-3-small (2.7 billion params) and GPT-3-medium (6.7 billion params) [Brown et al., 2020] achieve a 25% success rate — the exact same as random guessing. A large GPT-2 model (1.5 billion params) [Radford et al., 2018] achieves a slightly better 32.4% accuracy, but this is still only 7% higher than random prediction. Our NRGPT model at 125M parameters (an order of magnitude smaller than the MMLU results of GPT-2) achieves a 29.3% accuracy, which is the best performance per parameter count on this task that we are aware of.